OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. Solution:2, only solution

OpenStudy (anonymous):

(x-2)(x-2)=x^2-4x+4

OpenStudy (anonymous):

x-2=0 x=2